# Magneto-optical transport in tilted type-I double-Weyl semimetals in the presence of orbital magnetic moment

### v1

@@ Abstract
The magneto-optical transport of gapless type-I tilted-single Weyl semimetals exhibits suppression of total magnetoconductivities in the presence of orbital magnetic moment in linear and non-linear responses (Yang Gao et al, Phys. Rev. B 105, 165307 (2022)). In this work, we extend our study to investigate magnetoconductivities in gapless type-I tilted double-Weyl semimetals within semiclassical Boltzmann approach and show the differences that arise compared to single Weyl semimetals.
@@ I Introduction
Weyl semimetal is a three-dimensional topological state of matter, in which the conduction and valence bands touch at a finite number of nodes, called Weyl nodes[1, 2, 3, 4]. The Weyl nodes always appear in pairs due to the Nielsen-Ninomiya theorem[5, 5]. Each Weyl node can be regarded as a monopole in k-space carrying the topologcal charge n=1. Weyl semimetal has the Fermi arc surface states that connect the surface projections of two Weyl nodes [6, 7, 8, 9].
However, the topological charge of the Weyl node can be greater than one, namely $J>1$, and the corresponding materials are termed as multi-Weyl semimetals [10, 11, 12, 13]. For J=2, which is refered as double Weyl semimetal (DWSM), the dispersion relation in the vicinity of weyl node is quadratic in two symmetry directions and linear in the third direction. These Weyl nodes are protected by the crystallographic point group symmetries [10]. Density functional theory proposed DWSM in HgCr2Se4 [6] and SrSi2 [11] and can be achieved in photonic crystals [12]. Numerical studies show the presence of multiple surface Fermi arcs in multi-Weyl semimetal [14]. Multi-WSMs show some intriguing transport phenomena [13, 15, 16, 17, 18, 19, 20, 21]
In this paper, we study the linear and nonlinear magneto-optical responses for tilted DWSMs in the presence of an orbital magnetic moment. This has been studied for isotropic or single WSMs [22]. The orbital magnetic moment can be thought as self-rotation of the Bloch wave packet , and modify the energy of the Bloch electron under the external magnetic field [23]. This orbital moment change the magneto-optical responses of tilted Weyl semimetals [22]. We derive an analytic expression for the magnetoconductivity by means of the semiclassical Boltzmann approach. It is found that the orbital magnetic moment induces non-trivial magnetoconductivity term, which gives rise to a partial cancellation of the total magnetoconductivity. This cancellation is more pronounced compared to isotropic WSMs. Further, we analyzed this suppressed feature for linear (B-linear) and quadratic (quadratic-B) contribution in the magnetic field to magnetoconductivities. We also show that the B-linear (quadratic-B) magnetoconductivity exhibits a behavior that is dependent (independent) of the chirality of the Weyl node in both linear or nonlinear response regime as in the case of single WSMs[22].
@@ II Model Hamiltonian and Semiclassical Boltzmann approach
The non-interacting low energy effective Hamiltonian for tilted multi-Weyl semimetals is given by [15, 16, 17, 18, 19, 20, 21],
$\mathcal{H}_{J}$ $=$ $\hbar v_{t}k_{z}\mathsf{1}_{2\times 2}+\alpha_{J}\hbar[(\hat{k}_{-})^{J}\sigma_{+}+(\hat{k}_{+})^{J}\sigma_{-}]+\chi\hbar\mathit{v}_{F}\hat{k}_{z}\sigma_{z}$ (1)
$+$ $\hbar\mathit{v}_{F}t_{s}k_{z}\sigma_{0},$
where $\sigma_{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ and $\hat{k}_{\pm}=\hat{k}_{x}\pm i\hat{k}_{y}$, J represents monopole charge, $\mathit{v}_{F}$ is the effective velocity along $\hat{z}$ direction and $\alpha_{J}$ is the material dependent parameter, e.g. $\alpha_{1}$ and $\alpha_{2}$ are the Fermi velocity and inverse of the mass respectively for the isotropic and double WSMs. The energy dispersion for DWSMs is given by $\epsilon_{\bm{k}}^{s}=\hbar t_{s}\mathit{v}_{F}k_{z}+s\hbar\sqrt{\alpha_{n}^{2}(\rho_{2})^{2}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ with $\rho_{2}=(k_{x}^{2}+k_{y}^{2})$. We will use semiclassical Boltzmann equations in this study.
In the presence of a static magnetic field $\bm{B}$ and a time varying electric field $\bm{E}$, the semiclassical equations of motion are [23, 24]
@@
The first term on the right-hand side of Eq. (2) is $\bm{v}_{\bm{k}}^{s}=\frac{1}{\hbar}\bm{\nabla}_{p}\tilde{\varepsilon}_{\bm{k}}^{s}$, defined in terms of an effective band dispersion $\tilde{\varepsilon}_{s}(\bm{k})$. In topological metals such as WSMs, this quantity acquires a term due to the intrinsic orbital moment ,i.e., $\tilde{\epsilon_{\bm{k}}^{s}}=\epsilon_{\bm{k}}^{s}-\bm{m}_{\bm{k}}^{s}\cdot\bm{B}$, while $\bm{m}_{\bm{k}}^{s}$ is the orbital moment induced by the semiclassical “self-rotation” of the Bloch wave packet. The term $\Omega_{\bm{k}}^{s}$ is the Berry curvature [23, 24]
@@
The general exressions for Berry curvature and orbital magnetic moment for multi-WSMs are [25]
$\bm{\Omega}_{\bm{k}}^{s}=\pm\frac{s}{2}\frac{nv\alpha_{n}^{2}k_{\perp}^{2n-2}}{\beta_{\bm{k},s}^{3}}\{k_{x},k_{y},nk_{z}\}$ (6)
$\bm{m}_{\bm{k},s}^{\pm}=\frac{s}{2}\frac{env\alpha_{n}^{2}k_{\perp}^{2n-2}}{\hbar\beta_{\bm{k},s}^{2}}\{k_{x},k_{y},nk_{z}\}$ (7)
where $\beta_{\bm{k},s}=\sqrt{\alpha_{n}^{2}(\rho_{2})^{2}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ in the case of DWSMs.
@@
where the factor $D=1+\frac{e}{\hbar}(\bm{\Omega}_{\bm{k}}^{s}\cdot\bm{B})$ modifies the phase space volume.
@@
Here, $\tilde{f}^{s}(\tilde{\epsilon_{\bm{k}}^{s}})$ is the electron distribution function. where $\tau$ is the relaxation time originating from the scattering of electrons by phonons, impurities, electrons, and other lattice imperfections [26].
@@
Eq.(10) can be solved by expanding the distribution function as power series in the electric field as
@@ III Linear response of double WSMs
For linear electric field response, we retain only first two terms of Eq.(12) and substitute Eq.(3) in Eq.(10)
@@
We expand Eq.(15) up to second order in magnetic field and obtain
$\tilde{f}_{1}^{s}=\frac{\tau}{(1-i\omega\tau)}\biggl{[}e\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$
$-\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})^{2}(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$
$-\frac{e^{2}}{\hbar^{2}}\Omega_{\bm{k}}^{s}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$ (16)
From Eqs.(3) and Eq.(III), the expression for current density at time t is given by
$\bm{j}_{1}=-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{1}^{s}$
$-\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{0}^{s}$ (17)
@@
where $\sigma_{ab}(\omega)$ is the frequency dependent conductivity.
@@ III.1 Calculations of longitudinal conductivities components without magnetic field
Substituting Eq.(III) without B terms into the first term of Eq.(III), we get longitudinal conductivities components
@@
At T=0 k, $-\frac{\partial f_{0}^{s}}{\partial\varepsilon_{\bm{k}}^{s}}=\delta(\varepsilon_{\bm{k}}^{s}-\mu)$, we get
$\sigma_{zz}^{0}(\omega)$ $=$ $\frac{\hbar\pi\mathit{v}_{F}^{2}\sigma_{D}}{4t_{s}^{2}}\Bigl{[}1-\sqrt{1-t_{s}^{2}}\Bigr{]}$ (20)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}\frac{t_{s}}{1-t_{s}^{2}}+\frac{1}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (21)
where $\sigma_{D}=\frac{e^{2}\tau\mu^{2}}{(1-i\omega t)2\pi^{2}\mathit{v}_{F}\hbar^{5}}$ is Drude frequency complex conductivity. The $\sigma_{zz}^{0}(\omega)$ are modifed comppared to single WSM while $\sigma_{xx}^{0}(\omega),\sigma_{xx}^{0}(\omega)$ have the same form as in the case of single WSMs.

### v2

# Magneto-optical transport in tilted type-I multi-Weyl semimetals in the presence of orbital magnetic moment

@@ Abstract
The magneto-optical transport of gapless type-I tilted single Weyl semimetals exhibits suppression of total magnetoconductivities in the presence of orbital magnetic moment(OMM) in linear and nonlinear responses (Yang Gao et al., Phys. Rev. B 105, 165307 (2022)). In this work, we extend our study to investigate magnetoconductivities in gapless type-I tilted multi-Weyl semimetals(mWSMs) within the semiclassical Boltzmann approach and show the differences that arise compared to single-Weyl semimetals.
@@ I Introduction
Weyl semimetal is a three-dimensional topological state of matter, in which the conduction and valence bands touch at a finite number of nodes, called Weyl nodes[1, 2, 3, 4]. The Weyl nodes always appear in pairs due to the Nielsen-Ninomiya theorem[5, 5]. Each Weyl node can be regarded as a monopole in k-space carrying the topological charge n=1. Weyl semimetal has the Fermi arc surface states that connect the surface projections of two Weyl nodes [6, 7, 8, 9].
However, the topological charge of the Weyl node can be greater than one, namely $J>1$, and the corresponding materials are termed as multi-Weyl semimetals(mWSMs) [10, 11, 12, 13]. For J=2, which is referred to as double Weyl semimetal (DWSM), the dispersion relation in the vicinity of the Weyl node is quadratic in two symmetry directions and linear in the third direction. These Weyl nodes are protected by the crystallographic point group symmetries [10]. Density functional theory proposed DWSM in HgCr2Se4 [6] and SrSi2 [11] and can be achieved in photonic crystals [12]. Numerical studies show the presence of multiple surface Fermi arcs in multi-Weyl semimetal [14]. Multi-WSMs show some intriguing transport phenomena [13, 15, 16, 17, 18, 19, 20, 21]
In this paper, we study the linear and nonlinear magneto-optical responses for tilted mWSMs in the presence of an orbital magnetic moment. This has been studied for isotropic or single WSMs [22]. However, non-linear magnetoconductivity has not been discussed in the literature with the combined effects of both the tilting and orbital magnetic moment terms [23, 24, 25]. The orbital magnetic moment can be thought of as the self-rotation of the Bloch wave packet, and modifies the energy of the Bloch electron under the external magnetic field [26]. This orbital moment changes the magneto-optical responses of tilted mWSMs [22]. We derive an analytic expression for the magnetoconductivity employing the semiclassical Boltzmann approach. It is found that the orbital magnetic moment induces a non-trivial magnetoconductivity term, which gives rise to a partial cancellation of the total magnetoconductivity. This cancellation is more pronounced compared to isotropic WSMs. Further, we analyzed this suppressed feature for linear and quadratic contributions in the magnetic field to magnetoconductivities. We also show that the linear-B (quadratic-B) magnetoconductivity exhibits a behavior that is dependent (independent) of the chirality of the Weyl node in both linear or nonlinear response regimes, as in the case of single WSMs[22].
The paper is organized as follows: In Sec.II, we begin with the model of a 3D multi Weyl semimetal with a tilt in the z direction, and then the semiclassical equations of motion for the dynamics of the electron wave packet in the electric and magnetic fields are presented. In Sec. III, the B-linear and quadratic-B magnetoconductivities including the orbital magnetic moment are obtained in the linear response regime, and analyzed in detail. In Sec. IV, we study second harmonic generation, and give the second harmonic conductivity formula as well as the further analysis for this result. We end with conclusions in Sec. V.
@@ II Model Hamiltonian and Semiclassical Boltzmann approach
The non-interacting low-energy effective Hamiltonian for tilted multi-Weyl semimetals is given by [15, 16, 17, 18, 19, 20, 21],
$\mathcal{H}_{J}=\alpha_{J}\hbar[(k_{-})^{J}\sigma_{+}+(k_{+})^{J}\sigma_{-}]+\chi\hbar\mathit{v}_{F}k_{z}\sigma_{z}+\hbar\mathit{v}_{F}t_{s}k_{z}\sigma_{0}$
where $\sigma_{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ and $k_{\pm}=k_{x}\pm ik_{y}$, J represents monopole charge, $\mathit{v}_{F}$ is the effective velocity along $\hat{z}$ direction and $\alpha_{J}$ is the material dependent parameter, e.g. $\alpha_{1}$ and $\alpha_{2}$ are the Fermi velocity and inverse of the mass respectively for the isotropic and double WSMs. The energy dispersion for mWSMs is given by $\epsilon_{\bm{k}}^{s}=\hbar t_{s}\mathit{v}_{F}k_{z}+s\hbar\sqrt{\alpha_{J}^{2}k_{\perp}^{2J}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ with $k_{\perp}=\sqrt{k_{x}^{2}+k_{y}^{2}}$. We will use semiclassical Boltzmann equations in this study.
In the presence of a static magnetic field $\bm{B}$ and a time varying electric field $\bm{E}$, the semiclassical equations of motion at the location r and the wave-vector k in a given band are [26, 27]
@@
where -e is the electron charge. The first term on the right-hand side of Eq. (2) is $\bm{v}_{\bm{k}}^{s}=\frac{1}{\hbar}\bm{\nabla}_{p}\tilde{\varepsilon}_{\bm{k}}^{s}$, defined in terms of an effective band dispersion $\tilde{\varepsilon}_{s}(\bm{k})$. In topological metals such as WSMs, this quantity acquires a term due to the intrinsic orbital moment,i.e., $\tilde{\epsilon_{\bm{k}}^{s}}=\epsilon_{\bm{k}}^{s}-\bm{m}_{\bm{k}}^{s}\cdot\bm{B}$, while $\bm{m}_{\bm{k}}^{s}$ is the orbital moment induced by the semiclassical “self-rotation” of the Bloch wave packet. The term $\Omega_{\bm{k}}^{s}$ is the Berry curvature [26, 27]
@@
The general expressions for Berry curvature and orbital magnetic moment for multi-WSMs are [28]
$\bm{\Omega}_{\bm{k}}^{s}=\pm\frac{s}{2}\frac{J\mathit{v}_{F}\alpha_{J}^{2}k_{\perp}^{2J-2}}{\beta_{\bm{k},s}^{3}}\{k_{x},k_{y},Jk_{z}\}$ (6)
$\bm{m}_{\bm{k},s}^{\pm}=\frac{s}{2}\frac{eJ\mathit{v}_{F}\alpha_{J}^{2}k_{\perp}^{2J-2}}{\hbar\beta_{\bm{k},s}^{2}}\{k_{x},k_{y},Jk_{z}\}$ (7)
where $\beta_{\bm{k},s}=\sqrt{\alpha_{J}^{2}k_{\perp}^{2}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ in the case of mWSMs.
@@
where the factor $D=1+\frac{e}{\hbar}(\bm{\Omega}_{\bm{k}}^{s}\cdot\bm{B})$ modifies the phase space volume [29].
@@
[…]
@@
Eq.(10) can be solved by expanding the distribution function as a power series in the electric field as
@@
Equation(13) compute the conductivity components under the combined influence of external electric and magnetic fields.
@@ III Linear response of double WSMs
For linear electric field response, we retain only the first two terms of Eq.(12) and substitute Eq.(8) in Eq.(10)
@@
We expand Eq.(15) up to the second order in magnetic field and obtain
$\tilde{f}_{1}^{s}$ $=$ $\frac{\tau}{(1-i\omega\tau)}\biggl{[}e\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$ (16)
- $\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})^{2}(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$
- $\frac{e^{2}}{\hbar^{2}}\Omega_{\bm{k}}^{s}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$
From Eqs.(8) and Eq.(16), the expression for current density at time t is given by
$\bm{j}_{1}$ $=$ $-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{1}^{s}$ (17)
- $\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{0}^{s}$
@@
where $\sigma_{ab}(\omega)$ is the frequency dependent conductivity. It is known that the single contribution from the group velocity $\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}$ or the Berry curvature $\bm{\Omega}_{\bm{k}}^{s}$ form the conventional longitudinal or Hall conductivities. In the presence of the magnetic field, the conductivity $\sigma(\omega)$ consists of the coupling terms between the group velocity $\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}$ and the Berry curvature $\bm{\Omega}_{\bm{k}}^{s}$ besides the conventional ingredients [see Eq. (17)]. Their combined contributions are triggered by the external magnetic field, and play a crucial role in the electron transport.
@@ III.1 Calculations of longitudinal conductivities components without magnetic field
Substituting Eq.(16) without B terms into the first term of Eq.(17), we get longitudinal conductivities components
@@
Refer to caption  Refer to caption  Refer to caption
Figure 1: The dependence of the optical conductivity on the tilt $t_{+}$ at zero B field for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are taken as $v_{F}=4.13\times 10^{5}$ m / s, $\alpha_{2}=0.009m^{2}/s$, $\alpha_{3}=4.5\times 10^{-11}m^{3}/s$, $\mu=1meV$, and $\tau=10^{-13}s$.
At T=0 K, $\frac{\partial f_{0}^{s}}{\partial\varepsilon_{\bm{k}}^{s}}=\delta(\varepsilon_{\bm{k}}^{s}-\mu)$, we get
for J=1,
$\sigma_{zz}^{0}(\omega)$ $=$ $\frac{\mu^{2}}{\pi\hbar^{2}\mathit{v_{F}}}\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}-t_{s}-\frac{1}{2}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}$ (20)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=\frac{\mu^{2}}{\pi\hbar^{2}\mathit{v_{F}}}\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}\frac{t_{s}}{(1-t_{s}^{2})}$ (21)
$+$ $\frac{1}{2}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}\frac{1}{2}$
for J=2,
$\sigma_{zz}^{0}(\omega)$ $=$ $\frac{\mu\mathit{v_{F}}}{4\hbar\alpha_{2}}\frac{\sigma_{D}}{t_{s}^{2}}\Bigl{[}1-\sqrt{1-t_{s}^{2}}\Bigr{]}$ (22)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=2\frac{\mu^{2}\sigma_{D}}{\pi\mathit{v_{F}}\hbar^{2}{t_{s}^{3}}}\Bigl{[}\frac{t_{s}}{1-t_{s}^{2}}$ (23)
$+$ $\frac{1}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}\frac{1}{2}$
for J=3
$\sigma_{zz}^{o}(\omega)$ $=$ $\frac{\mu^{\frac{2}{3}}\mathit{v_{F}}\sigma_{D}}{\pi^{\frac{1}{2}}\hbar^{\frac{2}{3}}\alpha_{3}^{2/3}}\Bigl{[}\frac{\Gamma(7/6)((-1+t_{s})(1-t_{s}^{2})^{1/3}+(1+t_{s})^{2/3}{}_{2}F_{1}(\frac{2}{3},\frac{4}{3},\frac{5}{3},\frac{2t_{s}}{-1+t_{s}})))}{3^{1/2}t_{s}(1-t_{s}^{2})^{2/3}\Gamma(2/3)}\Bigr{]}$ (24)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=3\frac{\mu^{2}}{\pi\hbar^{2}\mathit{v_{F}}}\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}\frac{t_{s}}{(1-t_{s}^{2})}+\frac{1}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}\frac{1}{2}$ (25)
Refer to caption  Refer to caption  Refer to caption
Figure 2: The frequency dependence of optical conductivity at t=0.5 at zero B-field for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
where $\sigma_{D}=\frac{e^{2}\tau}{(1-i\omega\tau)2\pi\hbar}$ is Drude frequency complex conductivity. The $\sigma_{zz}^{0}(\omega)$ are modified compared to single WSM while $\sigma_{xx}^{0}(\omega)$ have the same form as in the case of single WSMs. Eqs.(20) and (21). In the limit $t\rightarrow$0, $\sigma_{xx}^{0}(\omega)=\sigma_{yy}^{0}(\omega)=J\frac{\mu^{2}}{3\pi\hbar^{2}v_{F}}\sigma_{D}$,J=1 $\sigma_{zz}^{0}(\omega)=\frac{\mu^{2}}{3\pi\hbar^{2}v_{F}}\sigma_{D}$, J=2 $\sigma_{zz}^{0}(\omega)=\frac{\mu v_{F}}{8\hbar}\sigma_{D}$ and J=3 $\sigma_{zz}^{0}(\omega)=\sqrt{\frac{3}{\pi}}\frac{\Gamma[7/6]}{\Gamma[2/3]}\frac{\mu^{2/3}v_{F}}{5\hbar^{2/3}\alpha_{3}^{2/3}}\sigma_{D}$. Note that $\sigma_{xx}^{0}(\omega)\propto J\mu^{2}$ while $\sigma_{xx}^{0}(\omega)\propto\frac{\mu^{2/J}}{\alpha_{J}^{2/J}\hbar^{2/J}}$ obeys the power-laws in chemical potential. We find from Eqs.(20) to (25), the conductivity components are even with respect to the tilt parameter $t_{s}$. Thus, the total conductivity elements will be twice the contribution from single Weyl nodes. Fig.(1) shows a comparison between different m-WSMs on the tilt parameter $t_{+}$ dependence of $\sigma^{(0)}_{xx}(\omega)$ and $\sigma^{(0)}_{zz}(\omega)$. In addition, the frequency dependence of the conductivity $\sigma_{aa}^{(0)}(\omega)(a=x,z)$ exhibit Drude-type behavior as shown in Fig.(2).
Refer to caption  Refer to caption  Refer to caption
Figure 3: The dependence of the optical conductivity on the tilt $t_{+}=0.5$. for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)